OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −2) and perpendicular to the line y = 4x + 2? Write your answer as a fraction. I wrote down y = 4x + 2 on a graph and the coordinates, but I'm not sure how to find the slope that will overlap the coordinates.

OpenStudy (anonymous):

Perpendicular lines have slopes that are negative reciprocals of one another.

OpenStudy (jdoe0001):

haileemackk, do you know what a 'negative reciprocal' is?

OpenStudy (anonymous):

i think it's where you change the sign, but if i change it from y = 4x + 2 to y = 4x - 2, it doesn't overlap (4, -2)

OpenStudy (anonymous):

slope of is y = 4x + 2 m=4 since it is perpendicular m=-1/4 after that let x=0 and y=0 to the equation y = 4x + 2